# AF Embeddability of the C*-Algebra of a Deaconu-Renault Groupoid

### v1

@@ 8.2. The crossed product $C(X)\rtimes_{\sigma}\mathbb{Z}$
Now we apply Theorem 1 to show that it generalises Pimsner’s theorem [45, Theorem 9] (see also [9, Theorem 11.5] by Brown). In contrast to these references, we do not assume that $X$ is a compact metric space. Instead, we fix $X$ to be a locally compact, Hausdorff, second countable, and totally disconnected space.
@@
The definition above is stated differently in [9]. Recall that our topological space is more specific, being totally disconnected. For our purposes, both definitions are equivalent when $X$ is assumed to be compact. Now we show that Theorem 1 can be applied to prove [45, Theorem 9] (or [9, Theorem 11.5]).
@@ 8.3. Topological graph algebras
[52, Theorem 6.7] if $E$ is a compact topological graph with no sinks, then the following are equivalent:
@@
Note that, for a Deaconu-Renault groupoid $\mathcal{G}$ corresponding to $\sigma:X\rightarrow X$, then $E=(E^{0},E^{1},s,r)\coloneqq(X,X,\sigma,\mathrm{id})$ is a topological graph with $C^{*}(E)\cong C^{*}(\mathcal{G})$ by [27, Proposition 10.9]. For Deaconu-Renault groupoids, we prove the equivalence of items $(i),(ii),(iii)$ and $(v)$ in Theorem 8.12.
@@
otherwise $X$ would have an $\varepsilon$-pseudoloop based at $x$. Let $\Omega=\bigcup_{i=0}^{\infty}\sigma^{i}(V)$. Then $\Omega$ is compact open because $X$ is compact and totally disconnected. By 8.4, we have $\sigma(\Omega)\subset\Omega$ and $\sigma(\Omega)\cap V=\emptyset$. So $\sigma(\Omega)\subsetneq\Omega$. Let $f=1_{\sigma(\Omega)}$, define $h=\sigma_{\ast}(f)-f$. Then, for all $y\in X$, we have

### v2

@@ 6.2. The crossed product $C(X)\rtimes_{\sigma}\mathbb{Z}$
Now we apply Theorem 2 to show that it generalises Pimsner’s theorem [26, Theorem 9] (see also [5, Theorem 11.5] by Brown). In contrast to these references, we do not assume that $X$ is a compact metric space. Instead, we fix $X$ to be a locally compact, Hausdorff, second countable, and totally disconnected space.
@@
The definition above is stated differently in [5]. Recall that our topological space is more specific, being totally disconnected. For our purposes, both definitions are equivalent when $X$ is assumed to be compact. Now we show that Theorem 2 can be applied to prove [26, Theorem 9] (or [5, Theorem 11.5]).
@@ 6.3. Topological graph algebras
[36, Theorem 6.7] if $E$ is a compact topological graph with no sinks, then the following are equivalent:
@@
Note that, for a Deaconu-Renault groupoid $\mathcal{G}$ corresponding to $\sigma:X\rightarrow X$, then $E=(E^{0},E^{1},s,r)\coloneqq(X,X,\sigma,\mathrm{id})$ is a topological graph with $C^{*}(E)\cong C^{*}(\mathcal{G})$ by [15, Proposition 10.9]. For Deaconu-Renault groupoids, we prove the equivalence of items $(i),(ii),(iii)$ and $(v)$ in Theorem 6.12.
@@
otherwise $X$ would have an $\varepsilon$-pseudoloop based at $x$. Let $\Omega=\bigcup_{i=0}^{\infty}\sigma^{i}(V)$. Then $\Omega$ is compact open because $X$ is compact and totally disconnected. By 6.4, we have $\sigma(\Omega)\subset\Omega$ and $\sigma(\Omega)\cap V=\emptyset$. So $\sigma(\Omega)\subsetneq\Omega$. Let $f=1_{\sigma(\Omega)}$, define $h=\sigma_{\ast}(f)-f$. Then, for all $y\in X$, we have
@@
6.4. Cuntz-Pimsner algebras
Cuntz-Pimsner algebras form a large class of C*-algebras that includes, among others, the C*-algebra associated with a Deaconu-Renault groupoid $\mathcal{G}$. By applying the techniques described in [33], one can show that $C^{*}(\mathcal{G})$ is a Cuntz-Pimsner algebra, for which Schafhauser characterised the AF embeddability in [39, Theorem C]. We refer to Chapter 8 of [29] for an introduction to Cuntz-Pimsner algebras.
Definition 6.16.
Let $G$ be a locally compact, Hausdorff, second countable, étale groupoid, and let $c:G\rightarrow\mathbb{Z}$ be a continuous cocycle. We say that $c$ is unperforated if, for every $g\in G$ such that $c(g)=n>0$, there exist composable $g_{1},g_{2},\dots,g_{n}$ such that each $c(g_{i})=1$ and $g=g_{1}g_{2}\dots g_{n}$.
Lemma 6.17.
The canonical cocycle $c:\mathcal{G}\rightarrow\mathbb{Z}$ on the Deaconu-Renault groupoid $\mathcal{G}$ is unperforated.
Proof.
Let $(x,k,y)\in\mathcal{G}$ such that $c(x,k,y)=k>0$. Then there exists an $m$ such that $\sigma^{m}+k(x)=\sigma^{m}(y)$. Since $\sigma$ is surjective, we can choose $z\in X$ with $\sigma^{k-1}(z)=y$. This implies that
$\sigma^{m}+k(x)=\sigma^{m}(y)=\sigma^{m+k-1}(z).$
This implies that $(x,1,z)\in\mathcal{G}$ and $c(x,1,z)=1$. Moreover, for $i=1,\dots,k-1$, we have $(\sigma^{i-1}(z),1,\sigma^{i}(z))\in\mathcal{G}$ and $c(\sigma^{i-1}(z),1,\sigma^{i}(z))=1$, where $\sigma^{0}(z)=z$. Hence,
$(x,k,y)=(x,1,z)(z,1,\sigma(z))\dots(\sigma^{i-3}(z),1,\sigma^{i-2}(z))(\sigma^{i-2}(z),1,y).$
Therefore, $c$ is unperforated. ∎
We will apply the following theorem by Rennie, Robertson, and Sims [33] to write $C^{*}(\mathcal{G})$ as a Cuntz-Pimsner algebra.
Theorem 6.18.
[33, Lemma 4 and Theorem 11] Suppose $G$ is a locally compact, Hausdorff, second countable, étale groupoid, and that $c:G\rightarrow\mathbb{Z}$ is an unperforated continuous cocycle. Let $X(G)$ be the completion in $C_{r}^{*}(G)$ of the set
$\{f\in C_{c}(G):\mathrm{supp}\hphantom{.}f\subset c^{-1}(1)\}.$
Then the inclusion $I_{0}:C_{c}(c^{-1}(0))\rightarrow C_{c}(G)$ extends to an embedding $I_{0}:C_{r}^{*}(c^{-1}(0))\rightarrow C_{r}^{*}(G)$ and the inclusion $I_{1}:C_{c}(c^{-1}(1))\rightarrow C_{c}(G)$ extends to a linear map $I_{1}:X(G)\rightarrow C_{r}^{*}(G)$. The set $X(G)$ is a C*-correspondence over $C_{r}^{*}(G)$ with module actions
$a\cdot\xi=I_{0}(a)\xi\hskip 15.0pt\text{and}\hskip 15.0pt\xi\cdot a=\xi I_{0}(a)\hskip 15.0pt\text{for $\xi\in X(G)$, $a\in C_{r}^{*}(c^{-1}(0))$.}$
Moreover, the pair $(I_{1},I_{0})$ is a Cuntz-Pimsner covariant representation of $X(G)$, and the integrated form $I_{1}\times I_{0}$ is an isomorphism of the Cuntz-Pimsner algebra $\mathcal{O}_{C^{*}_{r}(c^{-1}(0))}(X(G))$ onto $C_{r}^{*}(G)$.
Remark 6.19.
We make a change in notation, writing $\mathcal{O}_{C^{*}_{r}(c^{-1}(0))}(X(G))$ instead of $\mathcal{O}_{X(G)}$ to maintain consistency with the notation used in Schafhauser’s paper [39].
Now we write $C^{*}(\mathcal{G})$ as a Cuntz-Pimsner algebra.
Corollary 6.20.
The C*-algebra $C^{*}(\mathcal{G})$ is isomorphic to $\mathcal{O}_{C^{*}(c^{-1}(0))}(X(\mathcal{G}))$.
Proof.
Since the canonical cocyle $c:\mathcal{G}\rightarrow\mathbb{Z}$ is unperforated, Theorem 6.18 implies that $C^{*}_{r}(\mathcal{G})$ is isomorphic to $\mathcal{O}_{C_{r}^{*}(c^{-1}(0))}(X(\mathcal{G}))$. The groupoids $\mathcal{G}$ and $c^{-1}(0)$ are amenable, then $\mathcal{O}_{C_{r}^{*}(c^{-1}(0))}(X(\mathcal{G}))=\mathcal{O}_{C_{*}(c^{-1}(0))}(X(\mathcal{G}))$ and $C^{*}_{r}(\mathcal{G})=C^{*}(\mathcal{G})$. ∎
Proposition 6.21.
[39, Proposition 4.3] Suppose $A$ and $B$ are C*-algebras and $H$ is a row-finite $A\text{--}B$ C*-correspondence. If $T$ is an $A$-Ferdholm operator, then $T\otimes 1_{H}$ is a $B$-Fredholm operator and the map $T\mapsto T\otimes 1_{H}$ induces a positive group homomorphism $[H]:K_{0}(A)\rightarrow K_{0}(B)$.
The following theorem by Schafhauser characterises the AF embeddability for a large class of Cuntz-Pimsner algebras.
Theorem 6.22.
[39, Theorem C] Suppose $A$ is an AF algebra and $H$ is a separable C*-correspondence over $A$. Then the following are equivalent:
  1. (1)
    $\mathcal{O}_{A}(H)$ is AF embeddable,
  2. (2)
    $\mathcal{O}_{A}(H)$ is quasidiagonal,
  3. (3)
    $\mathcal{O}_{A}(H)$ is stably finite.
Moreover, if $H$ is row-finite and faithful, the above conditions are equivalent to
  1. (4)
    if $x\in K_{0}(A)$ and $[H](x)\leq x$, then $[H](x)=x$.
Note that, when $X$ is totally disconnected, the C*-algebra $C^{*}(c^{-1}(0))$ is AF. Hence, for the C*-algebra $C^{*}(\mathcal{G})=\mathcal{O}_{C^{*}(c^{-1}(0))}(X(\mathcal{G}))$, conditions (1)–(3) are equivalent. It is not clear to the author, whether $X(\mathcal{G})$ is row-finite or faithful.